# SWAN: SGD with Normalization and Whitening Enables Stateless LLM Training

### v2

@@ Abstract
Adaptive optimizers such as Adam (Kingma & Ba, 2015) have been central to the success of large language models. However, they often require maintaining optimizer states throughout training, which can result in memory requirements several times greater than the model footprint. This overhead imposes constraints on scalability and computational efficiency. Stochastic Gradient Descent (SGD), in contrast, is a stateless optimizer, as it does not track state variables during training. Consequently, it achieves optimal memory efficiency. However, its capability in LLM training is limited (Zhao et al., 2024b). In this work, we show that pre-processing SGD in a stateless manner can achieve the same performance as the Adam optimizer for LLM training, while drastically reduces the memory cost. Specifically, we propose to pre-process the instantaneous stochastic gradients using normalization and whitening. We show that normalization stabilizes gradient distributions, and whitening counteracts the local curvature of the loss landscape. This results in SWAN (SGD with Whitening And Normalization), a stochastic optimizer that eliminates the need to store any optimizer states. Empirically, SWAN has the same memory footprint as SGD, achieving $\approx 50\%$ reduction on total end-to-end memory compared to Adam. In language modeling tasks, SWAN demonstrates comparable or even better performance than Adam: when pre-training the LLaMA model with 350M and 1.3B parameters, SWAN achieves a 2× speedup by reaching the same evaluation perplexity using half as many tokens.
@@ 1 Introduction
Can LLMs be trained efficiently using stateless optimizers?
There is a growing body of research that has contributed to answering this question positively by developing novel optimizers that reduce the memory requirements associated with tracking internal state variables during the training of LLMs while achieving similar or even speedup boost performance compared to Adam. For instance, some methods rely solely on tracking the first moment of gradients (Xu et al., 2024a; Jordan et al., 2024), while others introduce an additional one-dimensional tracking variable on top of first moments (Zhang et al., 2024b; Zhao et al., 2024c). Alternatively, approaches focusing exclusively on pre-conditioner tracking have also been proposed (Pooladzandi & Li, 2024; Li, 2017). Another line of work focuses on using low-rank approximations to store the first and second moments, thereby reducing the memory cost associated with tracking optimizer states (Lialin et al., 2023; Hao et al., 2024; Zhao et al., 2024a). More recently, Zhu et al. (2024) introduced an almost stateless optimizer for efficient LLM training, which proposes to only track rank-1 state variables during training.
@@
Our optimizer consists of combining two well-known operators to pre-process the raw gradients: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$. $\mathtt{GradNorm}$ applies a row-wise standardization on the gradient matrix, while $\mathtt{GradWhitening}$ orthogonalizes the normalized gradient matrix. We show that these operators aims at stabilizing the stochasticity of gradient distributions during training, and neutralizing the local geometry of the loss landscape, respectively. When applied together, both operators enables SWAN to rely solely on the statistics of the current gradient matrices. This approach eliminates the need to track state variables, thereby matching the memory footprint of SGD. In addition to memory savings, SWAN also demonstrates significant computational efficiency: our empirical evaluations on pre-training LLaMA (Touvron et al., 2023a) models on the C4 dataset with multiple model sizes show consistently the same or better performance than Adam and other low-rank optimizers. Remarkably, at the 350M and 1.3B scale, our method achieves up to 2X faster convergence in terms of tokens seen compared to Adam. Our contributions are summarized below:
@@
      Robustness to ill-conditioned problems: on LLM pre-training, SWAN consistently and efficiently converges without learning rate warm-up (Section 6.1, Section 6.4).
@@
      Memory efficiency: SWAN achieves the memory footprint of SGD, that is $\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states when compared to Adam.
@@
      Computational efficiency: consistently achieves the same or better performance when compared to Adam across several scales and can significantly improve token-efficiency during training up to 2$\times$ at the 1B scale for the LLaMA model.
  • Theoretical analysis of SWAN. We show that (1) $\mathtt{GradNorm}$ can stabilize the heterogeneous covariance of LLM gradients, leveraging the redundancies in LLM gradient flows (Theorem 1 in Appendix B); and (2) $\mathtt{GradWhitening}$ can be derived as a non-diagonal second-order update under a specific structural assumption of the Hessian (Section 5.3). Additionally, we highlight that in the quadratic case, $\mathtt{GradWhitening}$ leads to convergence rates that are robust to the condition number of the local curvature (Theorem 2, Section B.1).
  • Fast convergence for LLM pretraining. Through experiments (Section 6) on LLM pretraining tasks, we demonstrate that SWAN not only reduces memory overhead but also achieves similar or even better performance compared to Adam. Notably, SWAN achieves convergence speed-ups of over 2$\times$ in terms of the number of tokens used for both 350M and 1.3B models (Figure 1).
(a) 350 M LLaMA model Pretraining
(b) 1.3 B LLaMA model Pretraining
@@
(d) Effective training throughput (same batch size)
Figure 1: SWAN performance preview on LLM pretraining. (a) and (b): On both 350M and 1.3B LLama architectures, SWAN achieves $>$ 2X speed-up vs Adam in terms of tokens seen. (c): memory footprint. We directly measure end-to-end memory under full-model training, with batch size = 1 sequence. This is different to the layer-wise training setting Zhao et al. (2024a). SWAN achieves near-SGD optimizer memory reduction ($\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states). (d) Training throughput analysis on training 1.3 B model on 4 $\times$ A100, under constant batch size = 130K tokens. We present two metrics: absolute throughput, measured by number of training tokens consumed per second; and effective throughput, which is absolute throughput adjusted by the token efficiency of optimizer relative to Adam. SWAN closely reproduces the absolute throughput of Adam, while improving the effective throughput.
@@ 2 Related works
Towards Stateless Optimizers for LLM Training.
Adaptive optimizers generally rely on tracking internal state variables to perform weight updates, which can substantially increase memory consumption when training large models. Several recent works have successfully managed to reduce the memory requirements associated with storing additional state variables for training LLMs. Muon (Jordan et al., 2024; Bernstein & Newhouse, 2024b), a newly proposed optimizer, has demonstrated strong acceleration and memory saving for LLM training by simplifying shampoo-like optimizers (Gupta et al., 2018; Anil et al., 2021; Shi et al., 2023; Wang et al., 2024; Vyas et al., 2024; Peirson et al., 2022; Lin et al., 2024) and requiring only the tracking of a first-moment estimate. SGD-Sal (Xu et al., 2024b) only stores the first moment with a learning rate estimated at the beginning of training. Sign-based methods such as (Chen et al., 2023) have also demonstrated success on training transformer-based models by only tracking first moments. There are also several works that aim to enhance the memory efficiency of Adam by reducing the memory cost associated with second moments. Adam-mini (Zhang et al., 2024b) significantly reduces memory usage by storing only scalar values for each parameter block, while Adalayer (Zhao et al., 2024c) retains only the scalar average of the second moment for each layer. Alternatively, PSGD (Pooladzandi & Li, 2024; Li, 2017) focuses on exclusively tracking a pre-conditioner, eliminating the need to track a first moment estimate. Finally,  (Zhu et al., 2024) proposes an nearly stateless optimizer to train LLMs, which only requires tracking rank-1 first and second moments. However, all the aforementioned optimizers still require the storage of state variables. In contrast, SWAN completely eliminates the need to store internal states for both the first and second moments by employing a combination of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ steps, which is discussed next.
@@
Gradient pre-processing is a common technique used to enhance performance of optimizers. Various pre-processing procedures have been proposed in the literature, such as signed gradient (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023; Kunstner et al., 2023), gradient clipping (Zhang et al., 2020), normalization (Zhang et al., 2020; You et al., 2019), and whitening (Yang & Laaksonen, 2008; Kingma & Ba, 2015; Hwang, 2024; Jordan et al., 2024; Bernstein & Newhouse, 2024c; a; Carlson et al., 2015). In this work, we particularly focus on normalization and whitening. We apply normalization row-wise on gradient matrices, similar to (You et al., 2019); together with gradient whitening under a specific structural assumption of the Fisher Information (Hwang, 2024; Martens et al., 2018), recovering the orthogonalization step used in (Jordan et al., 2024; Tuddenham et al., 2022). Our key result is that composing normalization and whitening on stochastic gradients is sufficient to enable the efficient training of LLMs in a completely stateless manner. Compared to Lamb (You et al., 2019), our normalization operation is applied on raw gradients instead of Adam states. Compared to Muon (Jordan et al., 2024), SWAN removes first-moment tracking and instead uses normalization. We show empirically that removing any one of these two pre-processing steps from SWAN results in significant performance degradation (Section 6.4).
@@ 4 The SWAN Optimizer: Preprocessing SGD with Normalization and Whitening
Figure 2: Illustration of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ operators. In $\mathtt{GradNorm}$ operator, we perform standardization across the output dimensions (columns), using statistics computed row-wise. In $\mathtt{GradWhitening}$ operator (illustration adapted from Huang et al. (2019)), we treat each column of the gradient matrix ${\mathbf{G}}$ as a separate data sample. Then, $\mathtt{GradWhitening}$ can be seen as stretching/squeezing the gradient data along all eigen directions, such that the covariance matrix is the identity matrix.
@@
     $\tilde{{\mathbf{G}}}^{(t)}\leftarrow\mathtt{GradNorm}({\mathbf{G}}^{(t)})$
@@
  Input: ${\mathbf{G}}^{m\times n}$ with $m\leq n$. Number of iterations $k$. Step size $\beta$ (default = 0.8).
@@
Algorithm 3 $\mathtt{GradNorm}$ Operator
  Input: ${\mathbf{G}}^{m\times n}$
  return  $\frac{{\mathbf{G}}-\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}\mathbf{1}_{n}^{\top}}{\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\frac{1}{n}\sum_{k=1}^{n}{\mathbf{G}}_{:,k})^{2}}\mathbf{1}_{n}^{\top}}$
Figure 3: SWAN Optimizer. Note that for $\mathtt{GradWhitening}$, in principle, any approximate algorithms can be applied, such as matrix Pade approximate Song et al. (2022). Here we only show one particular working example based on Newton-Schulz, which introduces minimal overheads.
@@ 4.1 SWAN Update Rules
Consider the gradient matrix ${\mathbf{G}}\in\mathbb{R}^{m\times n}$ with rows and columns corresponding to input and output dimensions, respectively, of some block of model parameters. Let $1\leq i\leq m$ represent the input indices and $1\leq j\leq n$ represent the output indices. Instead of performing element-wise EMA to stabilize and normalize the noisy gradients, as done in Adam, we propose to standardize across the output dimensions at each time step $t$, as follows:
$\mathtt{GradNorm}({\mathbf{G}}):=\frac{{\mathbf{G}}-\bar{{\bm{g}}}\mathbf{1}_{n}^{\top}}{{\bm{s}}\mathbf{1}_{n}^{\top}}$ (1)
where $\bar{{\bm{g}}}:=\frac{1}{n}\sum_{j=1}^{n}{\mathbf{G}}_{:,j}$ is the mean across output dimension; ${\bm{s}}:=\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}-\bar{{\bm{g}}})^{2}}$ is the standard deviation across dimension; $\bar{{\bm{g}}}$ and ${\bm{s}}$ are ${m}$-dimensional column vectors; $\mathbf{1}_{n}$ is a ${n}$-dimensional column vector of ones.
$\mathtt{GradNorm}$ is the forward pass operator LayerNorm (Ba et al., 2016) (without learnable parameters) applied on backward gradients. Similar operations have been considered in the literature to process gradients (You et al., 2019). $\mathtt{GradNorm}$ allows the optimizer to be invariant under the following transformations on the gradient matrix: i) matrix-wise re-scaling; ii) matrix-wise re-centering; iii) row-wise re-scaling; and iv) row-wise re-centering. Compared to the sign operation ($\text{sign}({\mathbf{G}})$) and matrix-wise normalization ($\frac{{\mathbf{G}}}{\|{\mathbf{G}}\|}$) used in signed descent, $\mathtt{GradNorm}$ preserves richer information of the gradient scaling while offering invariance properties. In Section 5 we will show that when applied to a simplified transformer architecture, $\mathtt{GradNorm}$ also helps stabilize the gradient distribution.
@@
As discussed in Section 3, Adam relies on a second moment estimate to perform element-wise gradient whitening. More formally, the second moment of Adam estimate a diagonal approximation of the Fisher information matrix (FIM) via (Kingma & Ba, 2015; Hwang, 2024):
@@
where the exponent $-\frac{1}{2}$ stands for matrix inverse square root. $({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}$ is simply the orthogonalization of ${\mathbf{G}}$, i.e., the closest orthogonal matrix to ${\mathbf{G}}$ (w.r.t the Frobenius norm). The derivation of Equation 2 as structured FIM/Hessian approximation is discussed in Section 5.3.
@@
4.2 Practical Considerations

### v3

@@ Abstract
Adaptive optimizers such as Adam (Kingma & Ba, 2015) have been central to the success of large language models. However, they often require maintaining optimizer states throughout training, which can result in memory requirements several times greater than the model footprint. This overhead imposes constraints on scalability and computational efficiency. Stochastic Gradient Descent (SGD), in contrast, is a stateless optimizer, as it does not track state variables during training. Consequently, it achieves optimal memory efficiency. However, its capability in LLM training is limited (Zhao et al., 2024b). In this work, we show that pre-processing SGD using normalization and whitening in a stateless manner can achieve the same performance as the Adam optimizer for LLM training, while maintaining the same memory footprint of SGD. Specifically, we show that normalization stabilizes gradient distributions, and whitening counteracts the local curvature of the loss landscape. This results in SWAN (SGD with Whitening And Normalization), a stochastic optimizer that eliminates the need to store any optimizer states. Empirically, SWAN achieves $\approx 50\%$ reduction on total end-to-end memory compared to Adam. In language modeling tasks, SWAN demonstrates comparable or even better performance than Adam: when pre-training the LLaMA model with 350M and 1.3B parameters, SWAN achieves a 2× speedup by reaching the same evaluation perplexity using half as many tokens.
@@ 1 Introduction
Can LLMs be trained efficiently, without storing any optimizer states?
There is a growing body of research that has contributed to answering this question positively by developing novel optimizers that reduce the memory requirements associated with tracking internal state variables during the training of LLMs while achieving similar or even speedup boost performance compared to Adam. For instance, some methods rely solely on tracking the first moment of gradients (Xu et al., 2024a; Jordan et al., 2024), while others introduce an additional one-dimensional tracking variable on top of first moments (Zhang et al., 2024b; Zhao et al., 2024c). Alternatively, approaches focusing exclusively on pre-conditioner tracking have also been proposed (Pooladzandi & Li, 2024; Li, 2017). Another line of work focuses on using low-rank approximations to store the first and second moments, thereby reducing the memory cost associated with tracking optimizer states (Lialin et al., 2023; Hao et al., 2024; Zhao et al., 2024a). Finally, the concurrent work of  (Zhu et al., 2024) proposed an approximate gradient scaling method, that tracks the channel-wise or tensor-wise surrogate learning rates, further improving the memory efficiency.
@@
Our optimizer consists of combining two well-known operators to pre-process the raw gradients: $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$. $\mathtt{GradNorm}$ applies a row-wise normalization on the gradient matrix, while $\mathtt{GradWhitening}$ orthogonalizes the normalized gradient matrix. We show that these operators aims at stabilizing the stochasticity of gradient distributions during training, and neutralizing the local geometry of the loss landscape, respectively. When applied together, both operators enables SWAN to rely solely on the statistics of the current gradient matrices. This approach eliminates the need to track state variables, thereby matching the memory footprint of SGD. In addition to memory savings, SWAN also demonstrates significant computational efficiency: our empirical evaluations on pre-training LLaMA (Touvron et al., 2023a) models on the C4 dataset with multiple model sizes show consistently the same or better performance than Adam and other low-rank optimizers. Remarkably, at the 350M and 1.3B scale, our method achieves up to 2X faster convergence in terms of tokens seen compared to Adam. Our contributions are summarized below:
@@
      Memory efficiency: SWAN only requires the memory footprint of SGD (w/o momentum), that is: $\approx 50\%$ reduction on total memory, and $\approx 100\%$ reduction on optimizer states compared to Adam. This can be further reduced with LOMO technique (Lv et al., 2023) (Figure 1 (c)).
@@
      Sample efficiency: through experiments (Section 6) on LLM pretraining tasks, we demonstrate that SWAN not only reduces memory overhead but also achieves similar or even better performance compared to Adam. Notably, SWAN achieves convergence speed-ups of over 2$\times$ in terms of the number of tokens used for both 350M and 1.3B models (Figure 1).
@@
      Computational efficiency: we propose a efficient scheme to accelerate the computation of SWAN (Section 4.2), achieving a similar throughput to Adam without the need for distributed computation (Section 6.2) required by recent optimizers such as shampoo Gupta et al. (2018).
    4. 4.
      Robustness to hyperparameters: the hyperparameters of SWAN are lazily tuned via heuristics; and they are shared across all model sizes. Even when compared to tuned-baselines obtained with learning rate sweeps, SWAN still delivers strong performances on LLM pretraining. This is a promising indication of applicability to real-world problems.
  • Theoretical consistency with LLM dynamnics. We show that (1) $\mathtt{GradNorm}$ can stabilize the heterogeneous covariance of LLM gradients, leveraging the redundancies in LLM gradient flows (Theorem 1 in Appendix C); and (2) $\mathtt{GradWhitening}$ can be derived as a non-diagonal second-order update under a specific structural assumption of the Hessian (Section 5.3). Additionally, we highlight that in the quadratic case, $\mathtt{GradWhitening}$ leads to convergence rates that are robust to the condition number of the local curvature (Theorem 2, Section C.1).
Table 1: The summary of memory consumption breakdown of optimizers, and the corresponding learning rate expressiveness. We assume all optimizers are applied on a 2D tensor of $m$ by $n$ ($m<n$). The optimizer state is defined as all intermediate variables maintained over time; while expressiveness is defined as the number of adaptive learning rates that a optimizer can model per tensor.
Adam SGD SGD-Sal Apollo (rank $r$) Apollo (mini) GaLore (rank $r$) SWAN
Optimizer States $\downarrow$ $2mn$ 00 $mn$ $2nr+2+(mr)$  [^1] The factor $mr$ comes from the storage of random projection tensors. It can be reduced by only storing the random seeds; however, this might cause challenges in a distributed setting. $2n+2+(m)$ $2nr+mr$ 00
Model weights $\downarrow$ $mn$ $mn$ $mn$ $mn$ $mn$ $mn$ $mn$
Expressiveness $\uparrow$ $mn$ $1$ blockwise $n$ $1$ $mn$ $mn$
(a) 350M model pretraining
(b) 1.3B model Pretraining
@@
(d) Throughput w/o distributed processing
Figure 1: SWAN Performance on LLM Pretraining. (a) and (b): On both 350M and 1.3B LLama architectures, SWAN achieves over 2X speed-up compared to Adam in terms of tokens seen. (c): Memory footprint. We measure end-to-end memory usage during full-model training with a batch size of 1 sequence. SWAN achieves nearly 100% reduction in optimizer states memory, and 50% total memory reduction (up to 70% when combined with per-layer training technique Zhao et al. (2024a); Lv et al. (2023)), and (d): Training throughput of a 1.3B model without distributed gradient pre-processing. With our acceleration scheme (denoteed by SWAN), we closely match the throughput of Adam without the need for distributed computation (Shi et al., 2023). The rightmost bar shows the effective throughput, adjusted for the token efficiency of the optimizer relative to Adam.
@@ 2 Related works
Towards Stateless LLM Training.
Adaptive optimizers generally rely on tracking internal state variables to perform weight updates, which can substantially increase memory consumption when training large models. Several recent works have successfully managed to reduce the memory requirements associated with storing additional state variables for training LLMs. Muon (Jordan et al., 2024; Bernstein & Newhouse, 2024b), a newly proposed optimizer, has demonstrated strong acceleration and memory saving for LLM training by simplifying shampoo-like optimizers (Gupta et al., 2018; Anil et al., 2021; Shi et al., 2023; Wang et al., 2024; Vyas et al., 2024; Peirson et al., 2022; Lin et al., 2024) and requiring only the tracking of a first-moment estimate. SGD-Sal (Xu et al., 2024b) only stores the first moment with a learning rate estimated at the beginning of training. Sign-based methods such as (Chen et al., 2023) have also demonstrated success on training transformer-based models by only tracking first moments. There are also several works that aim to enhance the memory efficiency of Adam by reducing the memory cost associated with second moments. Adam-mini (Zhang et al., 2024b) significantly reduces memory usage by storing only scalar values for each parameter block, while Adalayer (Zhao et al., 2024c) retains only the scalar average of the second moment for each layer. Alternatively, PSGD (Pooladzandi & Li, 2024; Li, 2017) focuses on exclusively tracking a pre-conditioner, eliminating the need to track a first moment estimate. Finally,  (Zhu et al., 2024) proposes a optimizer based on approximate gradient scaling to train LLMs, which requires tracking channel-wise or tensor-wise approximation to element-wise adaptive learning rates. However, all the aforementioned optimizers still require the storage of state variables. In contrast, SWAN completely eliminates the need to store internal states for both the first and second moments by employing a combination of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ steps, which is discussed next.
@@
Gradient pre-processing is a common technique to improve optimizer performance. Various pre-processing procedures have been proposed in the literature, such as signed gradient (Bernstein et al., 2018; Crawshaw et al., 2022; Chen et al., 2023; Kunstner et al., 2023), gradient clipping (Zhang et al., 2020), normalization (Zhang et al., 2020; You et al., 2019), and whitening (Yang & Laaksonen, 2008; Kingma & Ba, 2015; Hwang, 2024; Jordan et al., 2024; Bernstein & Newhouse, 2024c; a; Carlson et al., 2015). In this work, we focus on normalization and whitening. We apply normalization row-wise on gradient matrices, together with gradient whitening under a specific structural assumption of the Fisher Information (Hwang, 2024; Martens et al., 2018), recovering the orthogonalization step used in (Jordan et al., 2024; Tuddenham et al., 2022). Our key result is that composing normalization and whitening on stochastic gradients can be sufficient to enable the efficient training of LLMs in a stateless manner. In our setting, we show empirically that both removing any one of these two pre-processing steps from SWAN results in significant performance degradation (Section 6.4). Compared to Lamb (You et al., 2019), our normalization operation is applied on raw gradients row-wise; while LAMB applied layer-wise global normalization on Adam states instead. Compared to Muon (Jordan et al., 2024), SWAN removes first-moment tracking and instead uses row-wise normalization, making the optimizer fully statless. Moreover, our proposed efficient heuristic scheme for computing square root inverse requires $\mathcal{O}(m^{2})$ instead of $\mathcal{O}(m^{3})$ of the standard NS scheme (Song et al., 2022; Li et al., 2018; Huang et al., 2019).
@@ 4 The SWAN Optimizer: Preprocessing SGD with Normalization and Whitening
Figure 2: Illustration of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ operators. In $\mathtt{GradNorm}$ operator, we perform standardization across the output dimensions (columns), using statistics computed row-wise. In $\mathtt{GradWhitening}$ operator (illustration adapted from Huang et al. (2019)), we treat each column of the gradient matrix ${\mathbf{G}}$ as a separate data sample. Then, $\mathtt{GradWhitening}$ can be seen as stretching/squeezing the data such that the covariance matrix is the identity across all eigen directions.
@@
     $\tilde{{\mathbf{G}}}^{(t)}\leftarrow\mathtt{GradNorm}({\mathbf{G}}^{(t)})$ (Eq (1))
@@
  Input: ${\mathbf{G}}^{m\times n}$ with $m\leq n$. Number of iterations $K$. Step size $\beta$. Boolean diag indicating if use diagonal substitution scheme.
@@
     if diag then
        ${\mathbf{Y}}\leftarrow\beta{\mathbf{Y}}\text{Diag}(3{\mathbf{I}}-{\mathbf{Z}}\text{Diag}({\mathbf{Y}}))$,
        ${\mathbf{Z}}\leftarrow\beta(3{\mathbf{I}}-\text{Diag}({\mathbf{Z}}){\mathbf{Y}})\text{Diag}({\mathbf{Z}})$
     else
@@
Figure 3: SWAN Optimizer. When $\mathtt{diag}$ is set to $\mathtt{True}$, we cover the fast variant denoted by SWAN.
@@ 4.1 SWAN Update Rules
Consider the gradient matrix ${\mathbf{G}}\in\mathbb{R}^{m\times n}$ with rows and columns corresponding to input and output dimensions, respectively, of some block of model parameters. Let $1\leq i\leq m$ represent the input indices and $1\leq j\leq n$ represent the output indices. Instead of performing element-wise EMA to stabilize and normalize the noisy gradients, as done in Adam, we propose to standardize across the output dimensions at each time step $t$:
$\mathtt{GradNorm}({\mathbf{G}}):=\frac{{\mathbf{G}}}{{\bm{s}}\mathbf{1}_{n}^{\top}}$ (1)
where ${\bm{s}}:=\sqrt{\frac{1}{n}\sum_{j=1}^{n}({\mathbf{G}}_{:,j}})$ is ${m}$-dimensional the root mean square across dimension ( ${m}$-dimensional column vector); $\mathbf{1}_{n}$ is a ${n}$-dimensional column vector of ones.
$\mathtt{GradNorm}$ is the forward pass operator RMS-Norm (Zhang & Sennrich, 2019) applied on backward gradients. $\mathtt{GradNorm}$ allows the optimizer to be invariant under matrix-wise and row-wise rescaling. In Section 5 we show that this is the key to stabilizing the LLM gradient distribution. Meanwhile, compared to the sign operation ($\text{sign}({\mathbf{G}})$) and matrix-wise normalization ($\frac{{\mathbf{G}}}{\|{\mathbf{G}}\|}$) alternatives (Zhang et al., 2020; You et al., 2019), $\mathtt{GradNorm}$ preserves richer information of gradient scaling while offering invariance properties.
@@
As discussed in Section 3, Adam relies on a second moment estimate to perform element-wise gradient whitening. More formally, its second moment estimate a diagonal approximation to the Fisher information matrix (FIM):  (Kingma & Ba, 2015; Hwang, 2024):
@@
where the exponent $-\frac{1}{2}$ stands for matrix inverse square root. $({\mathbf{G}}{\mathbf{G}}^{\top})^{-\frac{1}{2}}{\mathbf{G}}$ is simply the orthogonalization of ${\mathbf{G}}$, i.e., the closest orthogonal matrix to ${\mathbf{G}}$ (w.r.t the Frobenius norm). The derivation of Equation 2 as structured FIM/Hessian approximation is discussed in Section 5.3. Apart from the FIM/Hessian approximation view, Equation 2 can also be interpreted as minimizing the condition number of ${\mathbf{G}}$ (defined as the ratio between the largest and smallest singular values). It can be shown that after whitening, its condition number achieves the minimum value which is 1.
@@
4.2 Practical Considerations: Fast and Stable Implementation
Fast $\mathtt{GradWhitening}$